Solve the non-linear system of equations. Give Exact Values:
x^2+y^2=25
3x^2+2y^2=59
: Solve the non-linear system of equations. Give Exact Values:
x^2+y^2=25
3x^2+2y^2=59

Answer by ankor@dixie-net.com(4675)

x^2+y^2=25
3x^2+2y^2=59
:
Multiply the 1st equation by 2 and subtract from the 2nd equation:
3x^2+2y^2=59
2x^2+2y^2=50
--------------------subtraction eliminates y^2
x^2 = 9
x = sqrt(9)
x = +/-3
:
Find y
3^2+y^2=25
9+y^2=25
y^2=25-9
y^2 = 16
y = sqrt(16)
y = +/-4
:
:
Check solutions in 2nd equation
3(9)+2(16)=59